# Equal contribution    Project lead    Corresponding author

### v3

@@ 1 Introduction
Time series data is a major modality in real-world dynamic systems and applications across various domains (Box et al., 2015; Zhang et al., 2024; Liang et al., 2024). Analyzing time series data is challenging due to its inherent complexity and distribution shifts, yet it is crucial for unlocking insights that enhance predictive analytics and decision-making. As a key task in high demand, time series forecasting has long been studied and is vital for driving various use cases in fields such as energy, climate, education, quantitative finance, and urban computing (Jin et al., 2023; Nie et al., 2024; Mao et al., 2024). Traditionally, forecasting has been performed in a task-specific, end-to-end manner using either statistical or deep learning models. Despite their competitive performance, the field has not converged on building unified, general-purpose forecasting models until recently, with the emergence of a few foundation models (FMs) for universal forecasting (Das et al., 2024; Woo et al., 2024; Ansari et al., 2024). Although promising, they are generally small in scale and have limited task-solving capabilities compared to domain-specific models, limiting their real-world impact when balancing forecasting precision against computational budget.
@@ 2 Related Work
Time Series Forecasting. Deep learning models have become powerful tools for time series forecasting over the past decade, which can be broadly categorized into two types: (1) univariate models, such as DeepState (Rangapuram et al., 2018), DeepAR (Salinas et al., 2020), and N-BEATS (Oreshkin et al., 2020), which focus on modeling individual time series, and (2) multivariate models, which include both transformer-based approaches (Wen et al., 2023; Zhou et al., 2021; Nie et al., 2023; Liu et al., 2024b; Wang et al., 2024c; Chen et al., 2024) and non-transformer models (Sen et al., 2019; Jin et al., 2022; Wang et al., 2024b; Hu et al., 2024; Qi et al., 2024), designed to handle multiple time series simultaneously. While these models achieve competitive in-domain performance (Wang et al., 2025), many are task-specific and fall short in generalizability when applied to cross-domain data in few-shot or zero-shot scenarios.
Large Time Series Models. Pre-training on large-scale sequence data has significantly advanced modality understanding in language and vision domains (Dong et al., 2019; Selva et al., 2023). Building on this progress, self-supervised learning has been extensively developed for time series (Zhang et al., 2024), employing masked reconstruction (Zerveas et al., 2021; Nie et al., 2023) or contrastive learning (Zhang et al., 2022; Yue et al., 2022; Yang et al., 2023). However, these methods are limited in both data and model scale, with many focused on in-domain learning and transfer. Recently, general pre-training of time series models on large-scale data has emerged, though still in its early stages with insufficient exploration into sparse solutions. We discuss the development more in Appendix A. Unlike these dense models, Time-MoE introduces a scalable and unified architecture for pre-training larger forecasting foundation models, which is also more capable while maintaining the same scale of activated parameters or computational budget as dense models.
@@ 3.2.1 Time-300B Dataset
Training time series foundation models require extensive, high-quality data. However, well-processed large-scale datasets are still relatively scarce. Recent advancements have facilitated the collection of numerous time series datasets from various sources (Godahewa et al., 2021; Ansari et al., 2024; Woo et al., 2024; Liu et al., 2024d; a). Nonetheless, data quality still remains a challenge, with prevalent issues such as missing values and invalid observations (Wang et al., 2024a) that can impair model performance and destabilize training. To mitigate these issues, we developed a streamlined data-cleaning pipeline (Appendix C) to filter and refine raw data, and constructed the largest open-access, high-quality time series data collection named Time-300B for foundation model pre-training. Time-300B is composed of a diverse range of publicly available datasets, spanning multiple domains such as energy, retail, healthcare, weather, finance, transportation, and web, as well as a portion of synthetic data to enhance the data quantity and diversity. Time-300B covers a wide spectrum of sampling frequencies from seconds to yearly intervals, and has over 300 billion time points after being processed by our data-cleaning pipeline, as summarized in Table 1.
@@ 3.2.3 Model Configurations and Training Details
[…]
@@ 4 Main Results
Time-MoE consistently outperforms state-of-the-art models by large margins across 6 well-established benchmarks and settings (Appendix B). To ensure a fair comparison, we adhered to the configurations from (Woo et al., 2024) for out-of-distribution forecasting and (Wu et al., 2023a) for in-distribution forecasting with a unified evaluation pipeline we developed. Specifically, we evaluate Time-MoE against 16 different baselines, representing state-of-the-art models in long-term forecasting. They are categorized into two groups: (1) zero-shot forecasting evaluation group, includes pre-trained foundation models such as Moirai (2024), TimesFM (2024), Moment (2024), and Chronos (2024); (2) in-distribution (full-shot) forecasting evaluation group, consists of up-to-date models such as iTransformer (2024b), TimeMixer (2024b), TimesNet (2023a), PatchTST (2023), Crossformer (2023), TiDE (2023), DLinear (2023),FEDformer (2022b). We also include addition comparisons with Timer  (2024d), TFT (2021), and N-BEATS (2020) in Appendix D.3.
@@ 5 Conclusion
In this paper, we introduced Time-MoE, a scalable and unified architecture for time series foundation models that leverages a sparse design with mixture-of-experts to enhance computational efficiency without compromising model capacity. Pre-trained on our newly introduced large-scale time series dataset, Time-300B, Time-MoE was scaled to 2.4 billion parameters, with 1.1 billion activated, demonstrating significant improvements in forecasting accuracy. Our results validate the scaling properties in time series forecasting, showing that Time-MoE consistently outperforms dense models with equivalent computational budgets across multiple benchmarks. With its ability to perform universal forecasting and superior performance in both zero-shot and fine-tuned scenarios, Time-MoE establishes itself as a state-of-the-art solution for real-world forecasting challenges. This work paves the way for future advancements in scaling and enhancing the efficiency of time series foundation models.
@@ Appendix C Processed Data Archive
BDG-2 Bull Energy H 464 501,832 Wang et al. (2023b)
BDG-2 Cockatoo Energy H 4 17032 Wang et al. (2023b)
Covid19 Energy Energy H 1 31,912 Wang et al. (2023b)
@@
GEF12 Energy H 20 788,280 Wang et al. (2023b)
GEF17 Energy H 8 140,352 Wang et al. (2023b)
BDG-2 Hog Energy H 152 365,304 Wang et al. (2023b)
@@
Spanish Energy H 1 35,064 Wang et al. (2023b)

### v4

@@ 1 Introduction
Time series data is a major modality in real-world dynamic systems and applications across various domains (Box et al., 2015; Zhang et al., 2024; Liang et al., 2024). Analyzing time series data is challenging due to its inherent complexity and distribution shifts, yet it is crucial for unlocking insights that enhance predictive analytics and decision-making. As a key task in high demand, time series forecasting has long been studied and is vital for driving various use cases in fields such as energy, climate, education, quantitative finance, cloud service, and urban computing, (Jin et al., 2023; Nie et al., 2024; Wang et al., 2023c; Mao et al., 2024). Traditionally, forecasting has been performed in a task-specific, end-to-end manner using either statistical or deep learning models. Despite their competitive performance, the field has not converged on building unified, general-purpose forecasting models until recently, with the emergence of a few foundation models (FMs) for universal forecasting (Das et al., 2024; Woo et al., 2024; Ansari et al., 2024). Although promising, they are generally small in scale and have limited task-solving capabilities compared to domain-specific models, limiting their real-world impact when balancing forecasting precision against computational budget.
@@ 2 Related Work
Time Series Forecasting. Deep learning models have become powerful tools for time series forecasting over the past decade, which can be broadly categorized into two types: (1) univariate models, such as DeepState (Rangapuram et al., 2018), DeepAR (Salinas et al., 2020), and N-BEATS (Oreshkin et al., 2020), which focus on modeling individual time series, and (2) multivariate models, which include both transformer-based approaches (Wen et al., 2023; Zhou et al., 2021; Nie et al., 2023; Liu et al., 2024b; Wang et al., 2024c; Chen et al., 2024; Wang et al., 2022) and non-transformer models (Sen et al., 2019; Jin et al., 2022; Wang et al., 2024b; Hu et al., 2024; Qi et al., 2024; Wang, 2024), designed to handle multiple time series simultaneously. While these models achieve competitive in-domain performance (Wang et al., 2025), many are task-specific and fall short in generalizability when applied to cross-domain data in few-shot or zero-shot scenarios.
Large Time Series Models. Self-supervised learning has been extensively developed for time series (Zhang et al., 2024), employing masked reconstruction (Zerveas et al., 2021; Nie et al., 2023) or contrastive learning (Zhang et al., 2022; Wang et al., 2023b; Yue et al., 2022). However, these methods are limited in both data and model scale, with many focused on in-domain learning and transfer. Recently, general pre-training of time series models on large-scale data has emerged (Liang et al., 2024), though still in its early stages with insufficient exploration into sparse solutions. See Appendix A for more information. Unlike these dense models, Time-MoE introduces a scalable, unified architecture for pre-training larger, more capable forecasting foundation models while maintaining the same scale of activated parameters and computational budget.
@@ 3.2.1 Time-300B Dataset
Training time series foundation models require extensive, high-quality data. Recent advancements have facilitated the collection of numerous time series datasets from various sources (Godahewa et al., 2021; Ansari et al., 2024; Woo et al., 2024; Liu et al., 2024d; a). Nonetheless, data quality still remains a challenge, with prevalent issues such as missing values and invalid observations (Wang et al., 2024a) that can significantly impair model performance and destabilize training. To mitigate these issues, we developed a streamlined data-cleaning pipeline (Appendix C) to filter and refine raw data, and constructed the largest open-access, high-quality time series data collection named Time-300B for foundation model pre-training. Time-300B comprises a diverse array of publicly available datasets from domains such as energy, retail, healthcare, weather, finance, transportation, and web, augmented with synthetic data to enhance both quantity and diversity. It spans sampling frequencies from seconds to yearly intervals and, after processing through our data-cleaning pipeline, includes over 300 billion time points, as summarized in Table 1.
@@ 3.2.3 Model Configurations and Training Details
Informed by the scaling laws demonstrated in (Dubey et al., 2024; Touvron et al., 2023), which show that a 7- or 8-billion parameter model continues to improve performance even after training on over one trillion tokens, we chose to scale Time-MoE up to 2.4 billion parameters with around 1 billion of them activated. This model, Time-MoEultra, supports inference on consumer-grade GPUs with less than 8GB of VRAM. We have also developed two smaller models: Time-MoEbase, with 50 million activated parameters, and Time-MoElarge, with 200 million activated parameters, both specifically designed for fast inference on CPU architectures. The detailed model configurations are in Table 2. Each model undergoes training for $100,000$ steps with a batch size of $1024$, where the maximum sequence length is capped at $4096$. This setup results in the consumption of $4$ million time points per iteration. We choose $\left\{1,8,32,64\right\}$ as different forecast horizons in the output projection and set the factor of the auxiliary loss $\alpha$ to $0.02$. Refer to Appendix B for optimization details.
@@ 4 Main Results
Time-MoE consistently outperforms state-of-the-art models by large margins across 6 well-established benchmarks and settings (Appendix B). To ensure a fair comparison, we adhered to the configurations from (Woo et al., 2024) for out-of-distribution forecasting and (Wu et al., 2023a) for in-distribution forecasting with a unified evaluation pipeline we developed. Specifically, we evaluate Time-MoE against 16 different baselines, representing state-of-the-art forecasting foundation models. They are categorized into two groups: (1) zero-shot forecasting group, includes pre-trained models such as Moirai (2024), TimesFM (2024), Moment (2024), and Chronos (2024); (2) in-distribution (full-shot) forecasting group, consists of up-to-date models such as iTransformer (2024b), TimeMixer (2024b), TimesNet (2023a), PatchTST (2023), Crossformer (2023), TiDE (2023), DLinear (2023),FEDformer (2022b). We also include addition comparisons with Timer  (2024d), TFT (2021), and N-BEATS (2020) in Appendix D.3.
@@ 5 Conclusion
In this paper, we introduced Time-MoE, a scalable and unified architecture for time series foundation models that leverages a sparse design with mixture-of-experts to enhance computational efficiency without compromising model capacity. Pre-trained on our newly introduced large-scale time series dataset, Time-300B, Time-MoE was scaled to 2.4 billion parameters, with 1.1 billion activated, demonstrating significant improvements in forecasting capabilities. Our results validate the scaling properties in time series forecasting, showing that Time-MoE consistently outperforms dense models with equivalent computational budgets across multiple widely accepted benchmarks. With its ability to perform universal forecasting and superior performance in both zero-shot and fine-tuned scenarios, Time-MoE establishes itself as a state-of-the-art solution for real-world forecasting challenges. This work lays the groundwork for future advancements in scaling and enhancing the efficiency of time series foundation models, paving the way toward time series general intelligence.
Acknowledgement
Y. Nie acknowledges financial support from Princeton Language and Intelligence at Princeton University. M. Jin was supported in part by the NVIDIA Academic Grant Program and CSIRO – National Science Foundation (US) AI Research Collaboration Program.
@@ Appendix C Processed Data Archive
BDG-2 Bull Energy H 464 501,832 Wang et al. (2023d)
BDG-2 Cockatoo Energy H 4 17032 Wang et al. (2023d)
Covid19 Energy Energy H 1 31,912 Wang et al. (2023d)
@@
GEF12 Energy H 20 788,280 Wang et al. (2023d)
GEF17 Energy H 8 140,352 Wang et al. (2023d)
BDG-2 Hog Energy H 152 365,304 Wang et al. (2023d)
@@
Spanish Energy H 1 35,064 Wang et al. (2023d)